# Will LLMs Scaling Hit the Wall? Breaking Barriers via Distributed Resources on Massive Edge Devices

### v1

@@ I.B. Derived from Larger Models
MINITRON (Muralidharan et al., 2024) 4B Transformer Distillation, Pruning 8T tokens from Nemotron-4
Orca/Orca 2 (Mitra et al., 2023; Mukherjee et al., 2023) 7B; 13B Transformer Distillation Orca 2 dataset, FLAN-v2 (Longpre et al., 2023)
MINIMA (Zhang et al., 2023a) 3B Transformer Distillation (from Llama-2-7B) Pile (Gao et al., 2020), Wudao
Dolly-v2 (Databricks, 2023) 3B; 7B Transformer Instruction tuning (from Pythia) Databricks-dolly-15k
LaMini-LM (Wu et al., 2024b) 61M-7B Transformer Distillation LaMini instruction dataset
@@ I.C. Model Compression Approaches
SparseGPT (Frantar et al., 2023) Various Transformer Unstructured Pruning Not applicable
Wanda (Sun et al., 2023) Various Transformer Unstructured Pruning Not applicable
LoRAPrune (Zhang et al., 2023b) Various Transformer Unstructured Pruning Not applicable
ShortGPT (Men et al., 2024) Various Transformer Structured Pruning Not applicable
BitNet/BitNet b1.58 (Wang et al., 2023; Ma et al., 2024) Various Transformer Quantization (QAT) Not applicable
QLoRA (Dettmers et al., 2023) Various Transformer Quantization, Low-Rank Various fine-tuning datasets
SqueezeLLM (Kim et al., 2023) Various Transformer Quantization (PTQ) Not applicable
@@ II. Alternative Architecture Models
Mamba (Gu & Dao, 2023) 125M-1.3B Mamba Pre-training Pile (Gao et al., 2020)
Rene (Wolf et al., 2024) 1.3B Mamba Pre-training Dolma-1.7 (AI2, 2023)
Zamba2 (Glorioso et al., 2024) 2.7B Mamba Pre-training Not specified
Hymba (Dong et al., 2024) 125M-1.5B Hymba Pre-training DCLM-Baseline (Li et al., 2024a)
xLSTM (Beck et al., 2024) 125M-1.3B xLSTM Pre-training SlimPajama (Systems & Computer, 2023)
RWKV (Peng et al., 2023) 169M-14B RNN Pre-training Pile (Gao et al., 2020)
Specialized FlanT5 (Fu et al., 2023) 250M-3B Encoder-Decoder Instruction Tuning GSM8K (Cobbe et al., 2021)
FlanT5 (Chung et al., 2022) 80M-3B Encoder-Decoder Instruction Tuning Muffin, T0-SF, SNI and CoT
T5 (Raffel et al., 2020) 60M-3B Encoder-Decoder Pre-training C4 (Raffel et al., 2020)
DistilBERT (Sanh et al., 2019) 66M Encoder-only Distillation (from BERT) Wikipedia, BookCorpus
TinyBERT (Jiao et al., 2020) 14.5M Encoder-only Distillation (from BERT) Wikipedia, BookCorpus
ALBERT (Lan et al., 2020) 12M-18M Encoder-only Pre-training (parameter sharing) Wikipedia, BookCorpus
@@ Appendix F Distributed Collaborative Frameworks
Distributed collaborative frameworks enable the deployment, training, and fine-tuning of language models across multiple devices or servers. Table 3 presents a comparison of prominent frameworks in this domain. These frameworks can be broadly categorized into three types: cloud-based platforms that offer centralized resources for distributed computing, federated learning systems that enable training across decentralized data sources while preserving privacy, and fully decentralized frameworks that distribute computation across peer nodes. Some frameworks like Neurosurgeon (Kang et al., 2017), MoE2 (Jin et al., 2025), Edgent (Li et al., 2018a), and Galaxy (Ye et al., 2024) focus on collaborative inference by partitioning models between edge devices and servers. Others, such as FedLLM (Wu et al., 2024c), FedFM (Chen et al., 2024b), FedPET (Li et al., 2024e), and Photon (Sani et al., 2024), specialize in federated fine-tuning of large language models while maintaining data privacy. These frameworks are essential for enabling efficient deployment of language models in resource-constrained environments and for scenarios requiring privacy preservation or operation in disconnected settings.
@@
exo-explore/exo (Labs, 2025) Decentralized MIT
Together AI (AI, 2023) Cloud Commercial
FLock Platform (FLock, 2023) Federated, Blockchain Apache 2.0
OpenDiloco (PrimeIntellect-ai, 2025) Decentralized Apache 2.0
FederatedScope (alibaba, 2024) Federated Apache 2.0
FedML (FedML-AI, 2024) Federated Apache 2.0
Flower (Flowerlab, 2025) Federated Apache 2.0
FATE-LLM (Fan et al., 2023b) Federated Apache 2.0
FedLLM (Ye et al., 2025) Federated CC BY-NC 4.0

### v2

@@ I.B. Derived from Larger Models
MINITRON [158] 4B Transformer Distillation, Pruning 8T tokens from Nemotron-4
Orca/Orca 2 [159, 160] 7B; 13B Transformer Distillation Orca 2 dataset, FLAN-v2 [161]
MINIMA [162] 3B Transformer Distillation (from Llama-2-7B) Pile [155], Wudao
Dolly-v2 [163] 3B; 7B Transformer Instruction tuning (from Pythia) Databricks-dolly-15k
LaMini-LM [164] 61M-7B Transformer Distillation LaMini instruction dataset
@@ I.C. Model Compression Approaches
SparseGPT [165] Various Transformer Unstructured Pruning Not applicable
Wanda [166] Various Transformer Unstructured Pruning Not applicable
LoRAPrune [167] Various Transformer Unstructured Pruning Not applicable
ShortGPT [168] Various Transformer Structured Pruning Not applicable
BitNet/BitNet b1.58 [169, 170] Various Transformer Quantization (QAT) Not applicable
QLoRA [171] Various Transformer Quantization, Low-Rank Various fine-tuning datasets
SqueezeLLM [172] Various Transformer Quantization (PTQ) Not applicable
@@ II. Alternative Architecture Models
Mamba [61] 125M-1.3B Mamba Pre-training Pile [155]
Rene [173] 1.3B Mamba Pre-training Dolma-1.7 [146]
Zamba2 [62] 2.7B Mamba Pre-training Not specified
Hymba [63] 125M-1.5B Hymba Pre-training DCLM-Baseline [136]
xLSTM [64] 125M-1.3B xLSTM Pre-training SlimPajama [65]
RWKV [66] 169M-14B RNN Pre-training Pile [155]
Specialized FlanT5 [71] 250M-3B Encoder-Decoder Instruction Tuning GSM8K [174]
FlanT5 [175] 80M-3B Encoder-Decoder Instruction Tuning Muffin, T0-SF, SNI and CoT
T5 [14] 60M-3B Encoder-Decoder Pre-training C4 [14]
DistilBERT [69] 66M Encoder-only Distillation (from BERT) Wikipedia, BookCorpus
TinyBERT [59] 14.5M Encoder-only Distillation (from BERT) Wikipedia, BookCorpus
ALBERT [60] 12M-18M Encoder-only Pre-training (parameter sharing) Wikipedia, BookCorpus
@@ Appendix F Distributed Collaborative Frameworks
Distributed collaborative frameworks enable the deployment, training, and fine-tuning of language models across multiple devices or servers. Table 3 presents a comparison of prominent frameworks in this domain. These frameworks can be broadly categorized into three types: cloud-based platforms that offer centralized resources for distributed computing, federated learning systems that enable training across decentralized data sources while preserving privacy, and fully decentralized frameworks that distribute computation across peer nodes. Some frameworks like Neurosurgeon [74], MoE2 [75], Edgent [76], and Galaxy [77] focus on collaborative inference by partitioning models between edge devices and servers. Others, such as FlexLoRA [83], FedFM [84], FedPETuning [85], and Photon [91], specialize in federated fine-tuning of large language models while maintaining data privacy. These frameworks are essential for enabling efficient deployment of language models in resource-constrained environments and for scenarios requiring privacy preservation or operation in disconnected settings.
@@
exo-explore/exo [73] Decentralized MIT
Together AI [176] Cloud Commercial
FLock Platform [177] Federated, Blockchain Apache 2.0
OpenDiloco [90] Decentralized Apache 2.0
FederatedScope [178] Federated Apache 2.0
FedML [179] Federated Apache 2.0
Flower [88] Federated Apache 2.0
FATE-LLM [89] Federated Apache 2.0
FedLLM [180] Federated CC BY-NC 4.0